# A Holistic Optimization Framework for Energy Efficient UAV-assisted Fog Computing: Attitude Control, Trajectory Planning and Task Assignment

### v2

@@ Abstract
Unmanned Aerial Vehicles (UAVs) have significantly enhanced fog computing by acting as both flexible computation platforms and communication mobile relays. In this paper, we propose a holistic framework that jointly optimizes the total latency and energy consumption for UAV-assisted fog computing in a three-dimensional spatial domain with varying terrain elevations and dynamic task generations. Our proposed framework considers three important and interdependent modules: attitude control, trajectory planning, and task assignment. We first establish a fuzzy proportional-integral-derivative control model to determine the UAV’s attitude. Then, we propose an enhanced Ant Colony System (ACS) based algorithm, that includes a safety value and a decoupling mechanism to overcome the convergence issue in classical ACS, to compute the optimal UAV trajectory. Finally, we design an algorithm based on the Particle Swarm Optimization technique, to determine where each offloaded task should be executed. Under our proposed framework, the outcome of one module would affect the decision-making in one other, providing a holistic perspective of the system and thus leading to improved solutions. We demonstrate by extensive simulation results that our proposed framework can significantly improve the overall performance, measured by latency and energy consumption, compared to existing baseline approaches.
Index Terms:
@@
I Introduction
@@
This paper considers a fog computing environment with a single UAVs deployed at the network edge, aiming at maximizing the energy efficiency by collaboratively controlling the attitudes and planning the trajectories, as well as determining the location for each computing task to be executed. Without loss of generality, we focus on quadrotor UAVs, which are considered ideal for computation offloading in IoT networks due to their enhanced flexibility [5, 6, 7, 8].
@@
Attitude control, trajectory planning, and task assignment policy in UAV-assisted fog computing are very important processes that significantly impact the overall efficiency of the system [9, 10, 11]. Specifically, attitude control (by adjusting pitch, roll, and yaw) ensures a stable and precise orientation of the UAV during operation, which is a necessary condition to maintain a high quality of communication with IoT devices and other fog nodes [12]. In addition, a stable attitude facilitates the UAV to effectively perform computation and storage tasks as a fog node. On the other hand, trajectory planning of the UAV can reduce its power consumption by identifying the most efficient path to collect data and tasks based on the locations of IoT devices [13]. Finally, task assignment refers to the process of deciding whether a specific task should be handled locally by the IoT device, processed in the fog layer (including the UAV), or offloaded to the central cloud. Specific objectives, including minimizing latency, maximizing throughput, or optimizing energy efficiency, can be achieved by smartly assigning tasks to appropriate devices. Task assignment is often jointly optimized with resource allocation, where transmission resources such as power and bandwidth are distributed among different transmission pairs in the network, to facilitate the transmission process and improve the overall efficiency. The four processes are inherently linked in UAV-assisted fog computing. For example, a UAV’s attitude control would ensure that it maintains optimal orientations while following a planned trajectory or processing a task. Also, when deciding whether to offload a certain task to the fog or the cloud and how to allocate relevant transmission resources, the energy consumption and latency for a certain UAV to reach the proximity of the IoT initiating the task along a planned trajectory should also be taken into account.
Existing studies have considered two or three processes for joint optimization. For example, Wei et al. [14] proposed three decision-making algorithms to solve the joint optimization problem involving energy consumption and mean delay. Our previous published conference paper [15] considered the joint optimization of trajectory planning and task assignment, in a fog computing system assisted by one single UAV. We proposed an optimization framework that integrated the classical Ant Colony System (ACS, also referred as Ant Colony Optimization/ACO in some existing studies) and Particle Swarm Optimization (PSO) algorithms. The framework would generate the optimal trajectory in a two-dimensional plane with randomly generated computing tasks and obstacles, and decide the optimal location (the IoT device, the UAV, or the central cloud) for each task to be executed.
In this work, we revise and extend the model in [15] to better align with the practical scenarios of UAV-assisted fog computing. Particularly, we consider a three-dimensional network space (compared to a two-dimensional plane as in [15]) with varying terrain elevations. We also add the attitude control and optimal structure design modules for a quadrotor UAV, to ensure the stability of flights, prevent potential structural damage, and reduce the latency when the UAV ascends or descends by two newly proposed Active Disturbance Rejection Control (ADRC) strategies [16]. Compared to most existing similar studies, (e.g.,[17, 18, 6, 7, 8]), which only considered the trajectory planning and flight attitude at independent static points, we jointly consider planning the optimal trajectory and determining the attitudes, taking into account the extra consumption required for changing attitude along the trajectory. Our research is expected to provide new useful insights in UAV-assisted fog computing applications for improving the overall performance and efficiency.
@@
  • We propose a holistic framework with multiple modules including attitude control, trajectory planning, resource allocation, and task assignment, to optimize the overall energy efficiency for UAV-assisted fog computing in a three-dimensional spatial domain with varying terrain elevations. To the best of our knowledge, this is the first work that jointly considers these modules together and takes advantage of their intricate interconnections. By integrating and coordinating the modules, our proposed framework is capable of achieving optimal performance and energy efficiency in the dynamic and complex environment of UAV-assisted fog computing.
@@
  • We establish a fuzzy proportional-integral-derivative control (PID) model for quadrator UAV attitude control. Compared to the classical PID control commonly used in existing studies, the fuzzy PID control can capture the dependencies among parameters and adapt to the dynamic environment. We demonstrate that the fuzzy PID control can effectively improve the stability in takeoff, cruise, and landing phases, and thus reduce both latency and energy consumption in UAV-assisted fog computing.
  • We propose a computationally efficient algorithm called ACS-DS (ACS with Decoupling and Safety values), to incorporate two mechanisms for decoupling and safety values in the ACS algorithm, to overcome the common issues of slow convergence and trapped in local optimum in classical ACS in UAV trajectory planning problems. We prove the convergence of the proposed algorithm and show that it has a polynomial time complexity. We also verify its superiority in terms of the convergence performance over classical ACS by numerical experiments.
@@
  • We propose a heuristic algorithm based on the idea of PSO to solve the joint optimization problem involving resource allocation and task assignment in UAV-assisted fog computing, given that an initial trajectory of the UAV has been determined. The heuristic algorithm is shown to be capable of overcoming the complexities in the non-convex optimization problem and obtaining the optimal task assignment and resource allocation decisions efficiently.
  • We demonstrate, through extensive numerical experiments, that our proposed holistic framework can reduce overall operational efficiency cost by more than $58\%$ compared to baseline implementations.
@@
The rest of this paper is organized as follows. Section II reviews recent advancements on attitude control, trajectory planning, and task assignment in UAV-assisted fog computing. Section III provides descriptions on the UAV architecture as well as key metrics at the system level. Section IV explains the formulation of the joint optimization problem. Section V describes the proposed computationally efficient algorithms to solve the problem in detail. Section VI demonstrates the improvements of the proposed algorithm by extensive numerical results. Section VII concludes the paper.
II Related Work
The fuzzy proportional-integral-derivative (PID) control system was adopted in attitude control and trajectory planning processes [19, 20]. It has been demonstrated that a proper combination of proportional-integral (PI) and proportional-derivative (PD) values and an optimally tuned fuzzy gain can better stabilize and control the attitude of quadrotor UAVs, as well as track their trajectories with smaller errors.
@@ II Related Work
For trajectory planning of UAVs, heuristic algorithms such as PSO [21], ACS [22], and genetic algorithm (GA) [23], have been adopted to overcome the space and computation complexities in such problems. Compared with another branch of approaches that use neural networks and deep learning as the key techniques (e.g., [24, 25, 26]), heuristic algorithms are more interpretable and less data dependent, and thus more appropriate for UAV-assisted fog computing scenarios where the operational environments are usually highly diverse and dynamic, and transparent heuristics are preferred for regulators to validate and verify the operation. Among the heuristic algorithms, the PSO [27] is one of the most commonly used techniques in such problems. However, one major concern of applying PSO in complex systems is that PSO may be converged prematurely to local optima [28].
@@
ACS, based on the study of ants searching for food, is another commonly adopted technique in discrete and continuous optimization problems [29, 30]. While ACS-based approaches are well-known for their robustness, they also suffer the disadvantage of being prone to local optima due to premature convergence like PSO. To overcome this issue, Ali et al. proposed the maximum-minimum ACS that improved the performance at the expense of slower convergence speed [31]. In addition, Wang et al. [32] presented an adaptive double-layer ant colony optimization algorithm (DL-ACS) based on an elitist strategy (ADAS) and an improved moving average algorithm (IMA) to solve a three-dimensional UAV trajectory planning problem. Liu et al. [33] also proposed an efficient online path-planning algorithm based on DL-ACS, that was shown to improve the optimal solution and convergence speed concurrently, and to enhance the stability of the algorithm.
@@
Meanwhile, resource allocation or task assignment problems have been extensively studied in existing studies. For example, Cui et al. [34] proposed a multi-agent Q-learning-based reinforcement learning (MARL) framework, where each agent independently executes the allocation algorithm to optimize the overall energy efficiency. Li et al. [35] considered a joint optimization problem involving trajectory optimization and task allocation, with the goal of minimizing UAV energy consumption and optimizing computation offloading and using successive convex approximation (SCA) technique to solve it. Wu et al. [36] proposed a cooperative multi-agent deep reinforcement learning framework, which combines task assignment and allocation of limited communication resources to minimize the overall energy consumption and delay.
@@
In trajectory planning and/or task assignment problems, one notable issue of applying ACS or PSO is that the algorithm may enter the deadlock state where one or more tasks wait endlessly for resources [37]. Zhou et al. [38] and Hou et al. [39] both proposed ACS-based algorithms with enhanced communication mechanisms to avoid deadlocks. While the proposed methods managed to reduce the likelihood of deadlocks, they did not eliminate the possibilities of such undesirable events. Another effort to overcome the deadlock problem is to establish a dead-corner table (or referred as a dead angle table[40], which updates the path strength using a penalty function to avoid the deadlock. While this approach effectively overcomes the deadlock problem, it introduces additional overheads in maintaining the table and thus is less efficient.
@@
Another popular emerging approach to solve UAV-related planning and optimization problem is the deep reinforcement learning (DRL) (e.g., [41, 42]). However, we focus on classical heuristic algorithms (such as ACS and PSO), as they are known to be more scalable, more adaptive and less computationally intensive compared to DRL-based approaches [43]. These advantages make them particularly well-suited for the dynamic and large-scale nature of UAV-assisted fog computing scenarios.
TABLE I: A summary and comparison of related studies
Research Process(es) focused Method(s) Objective(s)
[34] Trajectory Planning & Task Assignment MARL Max. long-term rewards
[35] Trajectory Planning & Task Assignment SCA Max. UAV energy efficiency
[9] Trajectory Planning & Task Assignment Convex Optimization Min. energy consumption and latency
[21] Task Assignment PSO Min. the overall workflow execution cost
[23] Task Assignment GA Min. communication costs
[14] Task Assignment Reinforcement learning Max. the platform’s utility
[31] Trajectory Planning MMACS Min. route length while avoiding collision
[28] Trajectory Planning Multi-Swarm PSO Max. the survivability of UAV
[33] Trajectory Planning Double-layer ACS Max. the distribution efficiency
[32] 3D-Trajectory Planning Double-layer ACS Max. Allocation Efficiency
[33] Trajectory Planning Double-layer ACS Max. Allocation Efficiency
[38],[39] Trajectory Planning Enhanced ACS Avoid path deadlock
[44],[40],[45] Trajectory Planning Enhanced ACS & Dead-corner table Avoid path deadlock & Min. Task transfer latency
[19] Attitude Control Traditional PID controller Control attitude
[20] Trajectory Planning & Attitude Control Fuzzy PID controller Control attitude and track trajectory
[36] Task Assignment Reinforcement learning Min. energy consumption and latency
@@
Our Work 3D Trajectory Planning & Task Assignment & Attitude Control PSO & Fuzzy control & ACS-DS Min. energy consumption and latency
@@
Our work in this paper combines the techniques mentioned above that have been demonstrated to be effective in respective processes, including fuzzy PID attitude control, ACS-based trajectory planning, and PSO-based task allocation and resource allocation. Following this foundation, we further improve several aspects of the entire framework, such as 1) the hardware of the quadrotor is designed to match the efficient performance of the fuzzy PID control system, 2) two enhanced anti-lockout mechanisms (decoupling and safety values) in the classical ACS for trajectory planning, and 3) a modified PSO approach with improved efficiency that is more appropriate for task assignment and resource allocation in large-scale UAV-assisted fog computing systems.
III System Model
III-A Network Structure and Components

### v3

# Energy-Aware Holistic Optimization in UAV-Assisted Fog Computing: Attitude, Trajectory, and Task Assignment

@@ Abstract
Unmanned Aerial Vehicles (UAVs) have significantly enhanced fog computing by acting as both flexible computation platforms and communication mobile relays. In this paper, we consider four important and interdependent modules: attitude control, trajectory planning, resource allocation, and task assignment, and propose a holistic framework that jointly optimizes the total latency and energy consumption for UAV-assisted fog computing in a three-dimensional spatial domain with varying terrain elevations and dynamic task generations. We first establish a fuzzy-enhanced adaptive reinforcement proportional-integral-derivative control model to control the attitude. Then, we propose an enhanced Ant Colony System (ACS) based algorithm, that includes a safety value and a decoupling mechanism to overcome the convergence issue in classical ACS, to compute the optimal UAV trajectory. Finally, we design an algorithm based on the Particle Swarm Optimization technique, to determine where each offloaded task should be executed. Under our proposed framework, the outcome of one module would affect the decision-making in another, providing a holistic perspective of the system and thus leading to improved solutions. We demonstrate by extensive simulation results that our proposed framework can significantly improve the overall performance, measured by latency and energy consumption, compared to existing mainstream approaches.
keywords:
@@
journal: Computer Networks
[…]
fundingfundingfootnotetext: This work was partly supported the Guangdong Provincial Key Laboratory of Interdisciplinary Research and Application for Data Science, BNU-HKBU United International College (Project code 2022B1212010006), and Guangdong Higher Education Upgrading Plan (2021-2025) UIC R0400001-22 and R0400024-22.conferenceconferencefootnotetext: A preliminary version of this paper was presented at IEEE HPCC 2022 in December 2022, and was included in its proceedings (DOI: 10.1109/hpcc-dss-smartcity-dependsys57074.2022.00217).
1 Introduction
@@
We consider a fog computing environment with a single UAV deployed at the network edge, aiming at maximizing the energy efficiency by collaboratively controlling the attitudes, planning the trajectories, allocating the transmission and computation resources, and assigning computing tasks to appropriate devices for execution. Specifically, attitude control (by adjusting pitch, roll, and yaw) ensures a stable and precise orientation of the UAV during operation, which is a necessary condition to maintain a high quality of communication with IoT devices and other fog nodes [27]. In addition, a stable attitude facilitates the UAV to effectively perform computation and storage tasks as a fog node. On the other hand, trajectory planning of the UAV can reduce its power consumption by identifying the most efficient path to collect data and tasks based on the locations of IoT devices [3]. Finally, task assignment refers to the process of deciding whether a specific task should be handled locally by the IoT device, processed in the fog layer (including the UAV), or offloaded to the central cloud based on real-time application-specific scenarios [30]. Specific objectives, including minimizing latency, maximizing throughput, or optimizing energy efficiency, can be achieved by assigning tasks to appropriate devices. Task assignment is often jointly optimized with resource allocation, where transmission resources such as power and bandwidth are distributed among different transmission pairs in the network, to facilitate the transmission process and improve the overall efficiency.
The four processes that we consider are inherently linked in UAV-assisted fog computing. For example, a UAV’s attitude control would ensure that it maintains optimal orientations while following a planned trajectory or processing a task. Also, when deciding whether to offload a certain task to the fog or the cloud and how to allocate relevant transmission resources, the energy consumption and latency for a certain UAV to reach the proximity of the IoT initiating the task along a planned trajectory should also be taken into account.
Existing studies have considered two or three processes for joint optimization. For example, Cheng et al. [5] proposed three decision-making algorithms to solve the joint optimization problem involving energy consumption and mean delay. Zhou et al. [49] proposed a two–time-scale optimization framework that jointly determines caching placement and task offloading decisions while adaptively adjusting the UAV trajectory. However, few considered all these aspects together in an interconnected manner. A summary of relevant studies is provided in Table LABEL:tab:related, and we will discuss them in more detail in Section 2.
This work substantially extends our earlier conference paper [24], to better capture the practical dynamics of UAV-assisted fog computing networks. The major improvements include: 1) While the previous study focused on a two-dimensional setting with simplified flight and communication assumptions, the present work develops a three-dimensional, terrain-aware network model that incorporates altitude variation and realistic environmental constraints; 2) We introduce an additional module (attitude control) for quadrotor UAVs, enabling the system to maintain stable flight, enhance link reliability under varying orientations, and mitigate latency during altitude transitions; and 3) We refine the algorithm for trajectory planning to better avoid deadlock by introducing decoupling and safety values mechanisms. These extensions allow the UAV to adapt its physical configuration and communication behavior in a coordinated manner, thereby improving overall energy efficiency, communication stability, and task execution performance beyond what was achieved in [24]. Compared to similar existing studies (e.g.,[35, 48]), which only considered the trajectory planning and flight attitude at independent static points, we jointly consider planning the optimal trajectory and determining the attitudes, taking into account the extra consumption required for changing attitude along the trajectory. Our research is expected to provide new useful insights in UAV-assisted fog computing applications for improving the overall performance and efficiency.
Although some of the individual strategies employed in this work, such as fuzzy, adaptive, and PID algorithms, are well-established, the novelty of our approach lies in how these algorithms are integrated and jointly optimized within a unified UAV-assisted fog computing framework. In contrast to most existing studies that treat control and networking processes separately, our framework co-designs attitude control, trajectory planning, task assignment, and resource allocation in a mutually dependent manner. This cross-layer integration enables the UAV to adapt its physical dynamics and communication decisions in real time according to network conditions, task demands, and environmental factors. As a result, the proposed approach improves system-level performance in terms of energy efficiency, communication reliability, and latency, thereby providing a new perspective on the joint optimization of control and communication functions in UAV-enabled fog computing networks.
@@
  • 1.
    From a computer and communication network perspective, this work proposes a unified optimization framework for UAV-assisted fog computing networks that jointly integrates attitude control, trajectory planning, task assignment, and resource allocation within a three-dimensional network topology. By coupling the UAV’s physical-layer control (attitude control) with network-layer decision making (resource allocation and task assignment), our cross-layer joint optimization framework captures the mutual dependencies among communication stability, mobility dynamics, and computational load distribution, thereby achieving adaptive connectivity, reduced latency, and improved overall network performance in realistic, terrain-aware IoT environments.
  • 2.
@@
    We develop a fuzzy-enhanced adaptive reinforcement proportional-integral-derivative (FEAR-PID) model for attitude management of quadrator UAVs. Compared with classical PID and conventional fuzzy-PID control commonly used in existing studies, FEAR-PID captures interdependencies among parameters and adapts dynamically to environmental changes. Our results demonstrate that FEAR-PID significantly enhances stability during takeoff, cruising, and landing phases, thereby effectively reducing latency and energy consumption in UAV-assisted fog computing systems.
  • 3.
    We propose a computationally efficient algorithm called ACS-DS (Ant Colony System with Decoupling and Safety values) for UAV trajectory planning. The ACS-DS algorithm integrates decoupling and safety value mechanisms to address typical limitations of classical ACS, such as slow convergence rates and susceptibility to local optima. Numerical experiments also confirm ACS-DS’s superior convergence performance relative to mainstream heuristic and reinforcement learning-based methods.
  • 4.
@@
    We propose a heuristic algorithm based on PSO principles to effectively resolve the resource allocation and task assignment problems, given that an initial trajectory of the UAV has been determined. The heuristic algorithm efficiently overcomes the inherent complexities of the underlying non-convex optimization problems, providing quasi-optimal solutions for task assignment and resource allocation decisions.
  • 5.
    We demonstrate, through extensive numerical experiments, that our proposed holistic framework can reduce overall operational efficiency cost by more than 67% reduction in overall operational costs compared to existing heuristic and reinforcement learning-based methodologies. Our analysis and results underscore the importance of considering interdependencies among attitude control, trajectory planning, resource allocation, and task assignment in UAV-assisted fog computing. Consequently, the holistic approach significantly outperforms methods optimizing individual components separately, highlighting the cumulative benefits of joint optimization in such environment.
@@
The rest of this paper is organized as follows. Section 2 reviews recent advancements on attitude control, trajectory planning, and task assignment in UAV-assisted fog computing. Section 3 provides descriptions on the UAV architecture as well as key metrics at the system level. Section 4 explains the formulation of the joint optimization problem. Section 5 describes the proposed computationally efficient algorithms to solve the problem in detail. Section 6 demonstrates the improvements of the proposed algorithm by extensive numerical results. Section 7 concludes the paper.
2 Related Work
2.1 Attitude Control
The proportional-integral-derivative (PID) control system is a fundamental approach widely used in the attitude control and trajectory planning of quadrotor UAVs [25, 7]. Building on this foundation, fuzzy PID control introduces adaptive capabilities by incorporating fuzzy logic to dynamically adjust the proportional-integral (PI) and proportional-derivative (PD) components, resulting in improved stability and reduced trajectory tracking errors  [46]. Recent studies further enhance these methods by leveraging reinforcement learning (RL). For instance, RL has been employed to regulate linear velocity while fuzzy logic manages angular velocity, achieving a complementary control strategy [41]. Another approach is the participation of RL in the construction of the PID controllers, which optimizes the gain parameters in real-time to achieve more accurate adaptive control under dynamic and complex environmental conditions [2].
2.2 Trajectory Planning
@@
For trajectory planning of UAVs, heuristic algorithms such as PSO [31], ACS [9, 44], and genetic algorithm (GA) [48], have been adopted to overcome the space and computation complexities in such problems. Compared with another branch of approaches that use neural networks and deep learning as the key techniques (e.g., [35, 36]), heuristic algorithms are more interpretable and less data dependent, and thus more appropriate for UAV-assisted fog computing scenarios where the operational environments are usually highly diverse and dynamic [23], and transparent heuristics are preferred for regulators to validate and verify the operation. Among the heuristic algorithms, the PSO [17] is one of the most commonly used techniques in such problems. However, one major concern of applying PSO in complex systems is that PSO may be converged prematurely to local optima [32].
@@
ACS, based on the study of ants searching for food, is another commonly adopted technique in discrete and continuous optimization problems [10, 34]. While ACS-based approaches are well-known for their robustness, they also suffer the disadvantage of being prone to local optima due to premature convergence like PSO. To overcome this issue, Wang et al. [37] presented an adaptive double-layer ant colony optimization algorithm (DL-ACS) based on an elitist strategy (ADAS) and an improved moving average algorithm (IMA) to solve a three-dimensional UAV trajectory planning problem. Recently, Yan et al. [44] proposed a chaotic-polarized-simulated ant colony optimization (CPS-ACO) algorithm, by incorporating chaotic mapping for initial pheromone distribution, a polarizing pheromone recording rule, and a simulated annealing mechanism, CPS-ACO demonstrated enhanced convergence speed and robustness against local optima in trajectory planning.
2.3 Task Assignment
@@
Resource allocation or task assignment problems have been extensively studied in existing studies. For example, Cui et al. [6] proposed a multi-agent Q-learning-based reinforcement learning (MARL) framework, where each agent independently executes the allocation algorithm to optimize the overall energy efficiency. Li et al. [21] both considered a joint optimization problem involving trajectory optimization and task allocation, with the goal of minimizing UAV energy consumption and optimizing computation offloading and using successive convex approximation (SCA) technique to solve it. Wu et al. [40] proposed a cooperative multi-agent deep reinforcement learning framework, which combines task assignment and allocation of limited communication resources to minimize the overall energy consumption and delay.
2.4 Joint Optimization
@@
One notable issue of applying ACS, GA or PSO in joint optimization involving trajectory planning and task assignment is that the algorithm may enter the deadlock state where one or more tasks wait endlessly for resources [11]. Hou et al. [13] proposed ACS-based algorithms with enhanced communication mechanisms to avoid deadlocks. While the proposed methods managed to reduce the likelihood of deadlocks, they did not eliminate the possibilities of such undesirable events. Another effort to overcome the deadlock problem is segmented planning and reintegration [22], which introduces self-healing mechanisms like In-Road Repairing and Intersection Repairing. These methods dynamically recalculate paths and penalize deadlocked routes using a negative reinforcement strategy. While this approach effectively avoids deadlocks and improves reliability, it incurs computational overhead due to frequent evaluations and path adjustments.
Another emerging approach to solving UAV-related planning and optimization problems is deep reinforcement learning (DRL) (e.g., [29, 38, 47]). Recently, Tan et al. [35] built a two-layer optimization framework by combining differential evolutionary algorithms and deep reinforcement learning to optimize the UAV deployment, tasks assignment and resource allocation efficiency when participating in multi-access edge computing. Proximal policy optimization-based deep reinforcement learning algorithms were used in [16] to learn the optimal offloading strategy for dependent tasks in a joint optimization problem that also involves trajectory planning. DRL-based approaches have been also jointly applied with other optimization modules in edge computing, such as blockchain consensus leader election and waiting time window decision [43], and estimation of resource requirements through digital twins [42].
@@
Compared with DRL-based approaches, the heuristic algorithms (such as ACS and PSO) are known to be more scalable, more adaptive and less computationally intensive [45]. These advantages make them particularly well-suited for the dynamic and large-scale nature of UAV-assisted fog computing scenarios. We will also compare their performance and computational efficiency in Section 6 of this paper.
Table 1: A SUMMARY AND COMPARISON OF RELATED STUDIES
Research Year Process(es) focused Method(s) Objective(s)
[8] 2022 TA Heuristic Harris Hawks Optimization Min. energy consumption
[31] 2022 TP Clustered-PSO Min. energy efficiency
[13] 2022 TP Enhanced ACS Avoid path deadlock
[22] 2022 TP Segmented ACS with Self-healing Routing Avoid path deadlock
[37] 2021 3D-TP Double-layer ACS Optimal trajectory
[48] 2024 3D-TP Genetic Algorithm (GA) & SCA Optimal trajectory
[44] 2024 3D-TP Chaotic-Polarized-Simulated scheme ACO Min. trajectory loss rate
[25] 2019 ATC Traditional PID controller Max. control accuracy
[46] 2024 ATC Fuzzy PID Controller Min. control bias
[2] 2024 ATC RL (Deep Q-Network) PID Controller Max. control accuracy
[40] 2023 TA & RA MARL & Stochastic Game Model Min. energy consumption and latency
[7] 2016 3D-TP & ATC Fuzzy PID controller Control attitude and track trajectory
[41] 2024 3D-TP & ATC RL PID Controller Avoidance of obstacles
[38] 2023 2D-TP & TA RL (Deep Q-Network) Min. energy consumption and latency
[6] 2020 3D-TP & TA MARL Max. long-term resource efficiency
[49] 2022 3D-TP & TA Successive convex approximation Min. latency
[16] 2023 3D-TP & TA DRL Min. latency
[39] 2019 2D-TP & TA & RA GA & Stepwise Approximation Min. energy consumption and latency
[21] 2020 3D-TP & TA & RA SCA & ADMM (Alternating Direction Method of Multipliers) Max. UAV energy efficiency
[29] 2023 3D-TP & TA & RA RL (MADDPG) Min. energy consumption and latency
[35] 2024 3D-TP & TA & RA Differential Evolutionary (DE) & RL (TD3) Optimal trajectory and convergence
Our Work 2025 3D-TP & TA & RA & ATC ACS-DS & PSO & FEAR-PID Min. energy consumption and latency
@@
Note: TP = Trajectory Planning, TA = Task Assignment, RA = Resource Allocation, ATC = Attitude Control
@@
Our work in this paper combines the techniques mentioned above that have been demonstrated to be effective in respective processes, including FEAR-PID attitude control, ACS-based trajectory planning, and PSO-based task allocation and resource allocation. Following this foundation, we further improve several aspects of the entire framework, such as 1) the hardware of the quadrotor is designed to match the efficient performance of the FEAR-PID control system, 2) two enhanced anti-lockout mechanisms (decoupling and safety values) in the classical ACS for trajectory planning, and 3) a modified PSO approach with improved efficiency that is more appropriate for task assignment and resource allocation in large-scale UAV-assisted fog computing systems.
3 System Model
3.1 Network Structure and Components